The arc advertising elasticity is 1.5 as advertising expenditures of Rs. 12 Lakhs, what will demand be at an advertising expenditure of Rs. 10 lakhs?
1
Expert's answer
2018-02-20T08:54:08-0500
A / PQ = - (Ea / Ep), where Ea / Ep = 1.5 and PQ changes from 12 to 10, that`s why A will be A = 10 * 1.5 = 15, but firstly it will be A =18, when advertising expenditures of Rs. 12 Lakhs. So demand will be 18